OpenStudy (anonymous):

8x+x(x-2)=0 f(x)=8x+x(x-2) - Find the x-intercepts

OpenStudy (anonymous):

start with \[f(x)=8x+x^2-2x=x^2+6x\]

OpenStudy (anonymous):

then if you want the x- intercepts (zeros) set \[x^2+6x=0\] \[x(x+6)=0\] \[x=0,\text{ or } x =-6\]